Select the best single answer.

The water-gas shift reaction plays a central role in the chemical methods for obtaining cleaner fuels from coal: CO(g) + H2O(g) ⇌ CO2(g) + H2(g)

At a given temperature, Kp = 2.7. If 0.13 mol of CO, 0.56 mol of H2O, 0.62 mol of CO2, and 0.47 mol of H2 are put in a 2.0−L flask, in which direction does the reaction proceed?

The reaction will proceed to the right OR the reaction is at equilibrium?
Answer #1

Given reaction is:

CO(g) + H2O(g) CO2(g) + H2(g)

Molarity = moles / Volume

[CO] = 0.13 mol/ 2.0 L = 0.065 M

[H2O] = 0.56 mol/ 2.0L = 0.28 M

[CO2] = 0.62 mol/ 2.0 L = 0.31 M

[H2] = 0.47 mol / 2.0 L = 0.235 M

Kc = ([CO2]*[H2]) / ( [H2O][CO]) = (0.31*0.235)/(0.065*0.28) = 4

Now, Moles of gaseous reactant = Moles of Gaseous product

so, Kp = Kc = 4 = Q(say)

But given Kp = 2.7 < Q

Reaction is not at equilibrium

So, reaction will shift backward to give more reactants i.e. toward left